Oneaboveall:

A parking garage charges $5.50 to park for 4 hours and $7.75 to park for 7 hours. If the cost is a linear function of the number of hours parked, what is the cost to park for 3 hours?

TheSmartOne:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @QuestionCoveBot All you have to do is 7.75-5.50. \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @QuestionCoveBot It's that easy. \(\color{#0cbb34}{\text{End of Quote}}\) Yeah, no it isn't mate.

ramen:

and you deleted ur response qcbot....

TheSmartOne:

We need to first make an equation to calculate the cost for parking for a certain number of hours. This is because, there might be a base charge and then an hourly charge for every hour you park after that. That's why it's not as simple as just subtracting the two numbers

TheSmartOne:

Let's say that y is the cost for parking, and that x is the number of hours parked we want to make the equation in the form of y = mx + b where m is the slope and b is our y-intercept We would have to first solve and get our m and our b, and then we can plug in x=3 to find the cost for parking for 3 hours b, our y-intercept, would be that initial base charge m, our slope, would be that additional constant charge that you get for every hour you park more

TheSmartOne:

so for 7 hours, we are charged 7.75 for 4 hours, we are charged 5.50 if we want to write it as a point, We said x was the number of hours parked And y is the total cost for parking Points are written as (x,y) So we have two points (4, 5.50) (7, 7.75) You have to first find the slope \(\sf slope = \frac{y_2-y_1}{x_2-x_1}\) Once you find the slope, just replace 'm' with that number in y = mx + b and then you can solve for b then You can plug in either point You can put x=4 and y=5.50 and then you can solve for b Or you can even do x=7 and y=7.75 and then solve for b Once you get b, go back to your equation and replace 'b' with that number in y = mx + b And at this point you should have a number for both m and b You then just plug in x = 3 Because we want to know the total cost (y) for when we're parked for 3 hours So just solve for y, and that's your final answer!!

TheSmartOne:

Let me know if you get stuck anywhere along the way or if you want me to check your answer. I'd be more than happy to assist you further :)